Three isotopes of each carbon and oxygen is taken then types of CO2 molecules which can be formed is
  1. 6
  2. 9
  3. 12
  4. 18

Grade:11

1 Answers

Sahani Kumar
5 years ago
Types of CO2 formed = (3 isotopes of carbon) * 2(3 isotopes of O2)
So types of CO2 formed = 3 *( 2*3)= 3*6 = 18 
Hence option (d)  is correct. 
Regards 
Askiitians members
If you like this answer please encourage us by approving this answer.